# FMARS: Annotating Remote Sensing Images for Disaster Management using Foundation Models

### v1

@@ 3.4 Experiments
Using the dataset annotated with our FMARS approach, we can train standard semantic segmentation models to evaluate the knowledge transfer ability to smaller and more deployable models. In our experiments, we adopt state-of-the art solutions based on Segformer [23]. To counteract the inherent inaccuracies in fully automated labeling and the consequently lower recall for categories such as high vegetation, we apply UDA techniques for improved stability during training. Specifically, we adopt DAFormer ([7] and Masked Image Consistency (MIC) [8], both based on self-training in a teacher-student framework paradigm. We select a separate full-size image from each event as our test set based on the average information content, for a total of 19 images, and we conduct a full training using pretrained ImageNet weights for the backbone components. To address the high precision and low recall of the generated labels, we train the models ignoring the background class and apply a confidence threshold to the Softmax outputs, selecting an optimal cutoff threshold $\tau$ for each model empirically (i.e., $\tau=0.75$ for DAFormer and $\tau=0.9$ for MIC). We conduct every experiment using a tile size of $512\times 512$ and random sampling, weighted by the entropy (i.e., information content) of the available label, for $30,000$ iterations using AdamW as optimizer. For the UDA components, we maintain their original configuration, except for the removal of the ImageNet feature distance. The obtained numerical results, in terms of accuracy and Intersection over Union (IoU) are displayed in Table 3. The scores highlight the challenge of the problem at hand, where the baseline solution, without any precautions, collapses during training. Both UDA solutions appear instead very effective, with MIC being the more robust solutions among the two. We note that every configuration displays apparent low accuracy and IoU values, particularly for the roads and high vegetation classes. However, we underline that these scores were computed against a robust albeit imperfect labels, and therefore represent an approximate result due to the lack of reliable ground truth annotations. This is evident in the qualitative results shown in Fig. 2, clearly demonstrate a high fidelity between the model predictions and the ground truth, with MIC exhibiting more robustness to the challenging roads and high vegetation classes. These findings highlight the effectiveness of FMARS as automated approach in leveraging foundation models for annotation tasks in remote sensing. With the necessary precautionary measures such as UDA techniques, the proposed pipeline allows the generation of accurate labels on a large scale, and even the knowledge transfer to smaller yet accurate downstream segmentation models in the absence of manually labeled datasets.
@@ 4 Conclusions
In this work, we propose FMARS, a pipeline for automated large-scale annotation of VHR imagery leveraging VFM. FMARS exploits the increasing availability of open-source images and the flexibility of promptable large models to automatically generate high-quality instance and semantic segmentation labels. As an example application, we focus on the critical domain of disaster management and construct the FMARS dataset, providing over 25 million annotations across 19 disaster events around the globe, derived from pre-event imagery from the Maxar Open Data initiative. As example application, we train state-of-the-art segmentation frameworks on the generated labels. Due to the potential inaccuracies in the automated labels, we employ UDA techniques to increase the robustness of the learned features to real-world scenarios. While the proposed approach demonstrates promising results, it has certain limitations.. First, the dataset currently focuses only on pre-event images. Second, the taxonomy is limited to three classes. Finally, achieving high recall requires significant computational effort, especially in downstream training. Future work may address these limitations by developing a more robust automated annotation pipeline, or by exploring zero-shot learning approaches for open-set or panoptic segmentation tasks.

### v2

@@ 3.4 Experiments
[…]
Table 4: Performance comparison of DAFormer and MIC on a manually labelled Maxar partition across different classes.
@@
Using the dataset annotated with our FMARS approach, we can train standard semantic segmentation models to evaluate the knowledge transfer ability to smaller and more deployable models. In our experiments, we adopt state-of-the art solutions based on Segformer [23]. To counteract the inherent inaccuracies in fully automated labeling and the consequently lower recall for categories such as high vegetation, we apply UDA techniques for improved stability during training. Specifically, we adopt DAFormer ([7] and Masked Image Consistency (MIC) [8], both based on self-training in a teacher-student framework paradigm. We select a separate full-size image from each event as our test set based on the average information content, for a total of 19 images, and we conduct a full training using pretrained ImageNet weights for the backbone components. To address the high precision and lower recall of the generated labels, as well as missing categories, we train the models in an open set context, ignoring the background class [24]. As simple baseline, we apply a confidence threshold to the Softmax outputs, empirically evaluating the optimal cutoff threshold $\tau=0.9$ for both models. We conduct every experiment using a tile size of $512\times 512$ and random sampling, weighted by the entropy (i.e., information content) of the available label, for $30,000$ iterations using AdamW as optimizer. For the UDA components, we maintain their original configuration, except for the removal of the ImageNet feature distance.
Given the automated pipeline and the low reliability of the obtained labels for performance measurement, we validate results against the FMARS labels, as well as a small sample of 45 manually labelled tiles, derived from crops of each image in the test set. Tables 3 and 4 present the numerical results in terms of accuracy and Intersection over Union (IoU), evaluated against the FMARS test set and manual labels, respectively. The scores highlight the challenge of the problem at hand, where the baseline solution, without any precautions, collapses during training. Both UDA solutions appear very effective, with DAFormer even surpassing the original FMARS labels on the manually produced ground truth. This is evident in the qualitative results shown in Fig. 2 that demonstrate a high fidelity between the model predictions and the ground truth, with DAFormer exhibiting more robustness to the challenging high vegetation class. These findings highlight the effectiveness of FMARS as an automated technique to leverage foundation models for annotation tasks in remote sensing. With the necessary precautionary measures such as UDA techniques, the proposed pipeline allows the generation of accurate labels on a large scale, and even the knowledge transfer to smaller yet accurate downstream segmentation models in the absence of manually labeled datasets.
@@ 4 Conclusions
In this work, we propose FMARS, a pipeline for automated large-scale annotation of VHR imagery leveraging VFM. FMARS exploits the increasing availability of open-source images and the flexibility of promptable large models to automatically generate high-quality instance and semantic segmentation labels. As an example, we focus on the critical domain of disaster management and construct the FMARS dataset, providing over 25 million annotations across 19 disaster events around the globe, derived from pre-event imagery from the Maxar Open Data initiative. As representative downstream application, we train state-of-the-art segmentation frameworks on the generated labels. Due to the potential inaccuracies in the automated labels, we employ UDA techniques to increase the robustness of the learned features to real-world scenarios. While the proposed approach demonstrates promising results, it has certain limitations. First, the dataset currently focuses only on pre-event images. Second, we limited the taxonomy to three classes in these tests. Finally, achieving high recall requires significant computational effort, especially in downstream training. Future work may address these limitations by developing a more robust automated annotation pipeline, or by exploring zero-shot learning approaches for open-set or panoptic segmentation tasks, with a potentially boundless taxonomy.